OpenStudy (anonymous):

Determine whether the ordered pair is a solution of the inequality (5,11); 4y - 5x > 8

OpenStudy (callisto):

Put x=5, y=11 into 4y - 5x , what is the value you get?

OpenStudy (callisto):

@lovejones2012

OpenStudy (anonymous):

I am not understanding

OpenStudy (anonymous):

Is it an inequality or solution?

OpenStudy (callisto):

To test if it is a solution, you need to check if the inequality is valid at the point. So, Put (5,11) into 4y - 5x . If the value you get is > 8, the order pair is a solution. If not, then it is not. So, the first thing you need to do is to put x=5 and y=11 into 4y-5x

OpenStudy (callisto):

What is the value you get for that?

OpenStudy (anonymous):

8?

OpenStudy (callisto):

4(11)-5(5) =?

OpenStudy (anonymous):

19

OpenStudy (callisto):

So, is 19 greater than 8?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

So it is not a solution right? Since its great than 8?

OpenStudy (anonymous):

Sorry it is a solution

OpenStudy (callisto):

Yes :) It is a solution

OpenStudy (anonymous):

:)

OpenStudy (callisto):

Do you know how to do it now?

OpenStudy (anonymous):

Thanks

OpenStudy (callisto):

Welcome

OpenStudy (anonymous):

How many teams recieve 2 points when they win and 1 point when they tie. One season, a team one 67 points. They won 11 more games than they tied. How many wins and how many ties did the team have?

OpenStudy (callisto):

let x be the no. of win and y be the no. of tie 2x+1y =67 x = 11+y Can you solve them?